OpenStudy (anonymous):

375 students went to Six flags over Georgia. 15 buses were filled and 15 students rode with their parents. How many students were on each bus? A) 20 b)34 c)24 D)30

OpenStudy (anonymous):

@Ashley1997 helppp @terenzreignz @harharf

OpenStudy (anonymous):

375-15=?

terenzreignz (terenzreignz):

15 students riding with their parents basically means 15 less students to take the buses

OpenStudy (anonymous):

@terenzreignz so which one of those options is right?

OpenStudy (anonymous):

375-15=360 360/15=24 So the answer is C

terenzreignz (terenzreignz):

That leaves 360 students to take the bus, so with 15 buses, 360 divided by 15 is 24